OpenStudy (jake124):

(7 + 2)^2 - 2^3 = this mean exponet (^2 ^3

OpenStudy (jake124):

@Luigi0210

jhonyy9 (jhonyy9):

7+2= ?

jhonyy9 (jhonyy9):

step by step - ok ?

OpenStudy (jake124):

9

jhonyy9 (jhonyy9):

yes right 9^2 = ?

OpenStudy (jake124):

81

jhonyy9 (jhonyy9):

9^2 = 9*9 = 81 so is right

jhonyy9 (jhonyy9):

2^3 = ?

OpenStudy (jake124):

8

jhonyy9 (jhonyy9):

yes so 81-8 = ?

OpenStudy (jake124):

73

jhonyy9 (jhonyy9):

very nice job this is the right answer ok. ?

OpenStudy (jake124):

73 is the answer

jhonyy9 (jhonyy9):

yes sure exactly

jhonyy9 (jhonyy9):

you ve solved it step by step ok. ?

OpenStudy (jake124):

thanks

jhonyy9 (jhonyy9):

np was my pleasure and anytime good luck bye bye

jhonyy9 (jhonyy9):

hope you understood it all right - yes ?

OpenStudy (jake124):

yes i do

jhonyy9 (jhonyy9):

ok great bye bye